A bond offers a coupon rate of 9%, paid annually, and has a maturity of 15 years. The current market yield is 7%. Face value is $1,000. If market conditions remain unchanged, what should be the Capital Gains Yield of the bond?
Answer #1
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =15
Bond Price =∑ [(9*1000/100)/(1 + 7/100)^k]     +   1000/(1 + 7/100)^15
                   k=1
Bond Price = 1182.16
Using Calculator: press buttons "2ND"+"FV" then assign
PMT = Par value * coupon %=1000*9/(100)
I/Y =7
N =15
FV =1000
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(7/(100),15,-9*1000/(100),-1000,)
capital gains yield = ((Ending price)/Beginning price-1)
=((1174.91)/1182.16-1)
=-0.61%
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =14
Bond Price =∑ [(9*1000/100)/(1 + 7/100)^k]     +   1000/(1 + 7/100)^14
                   k=1
Bond Price = 1174.91
Using Calculator: press buttons "2ND"+"FV" then assign
PMT = Par value * coupon %=1000*9/(100)
I/Y =7
N =14
FV =1000
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(7/(100),14,-9*1000/(100),-1000,)